OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1).

OpenStudy (anonymous):

slope intercept form is y=mx+C where m=slope=-3 put x =1 and y=1 to get constant C 1=-3(1)+C C=4 so equation of line is becomes Y=-3x+4

OpenStudy (anonymous):

thanks! @sami-21